A copper rod whose length is 10.0 is heated to 50°c. Find its New length take linear expansivity of copper as 0.000017


1
Expert's answer
2021-01-28T14:01:35-0500

The equation of linear heat expansion says

"l=l_0(1+\alpha \Delta t)\\
l=\rm 10.0\times (1+0.000017\times 50)= 10.0085\: m"